# Position: Many generalization measures for deep learning are fragile

### v2

@@ Abstract
A wide variety of generalization measures have been applied to deep neural networks (DNNs). Although obtaining tight bounds remains challenging, such measures are often assumed to reproduce qualitative generalization trends. In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks— are fragile: small training modifications that barely affect the underlying DNN can substantially change a measure’s value, trend, or scaling behavior. For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants can reverse the slope of a learning curve in widely used generalization measures like the path norm. We also identify subtler forms of fragility. For instance, the PAC-Bayes origin measure is regarded as one of the most reliable, and is indeed less sensitive to hyperparameter tweaks than many other measures. However, it completely fails to capture differences in data complexity across learning curves. This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behavior, in learning curves, but which is not a post-mortem measure. Beyond demonstrating that many bounds—such as path, spectral and Frobenius norms, flatness proxies, and deterministic PAC-Bayes surrogates—are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
@@ 1 Introduction
Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning. They were developed with two intertwined aims: to furnish tight mathematical guarantees and to explain generalization. For deep networks the same aspiration has driven a profusion of bounds. While tight bounds have proven elusive, with some exceptions (Dziugaite & Roy, 2017; Pérez‐Ortiz et al., 2020; Valle-Pérez & Louis, 2020; Lotfi et al., 2022), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs). These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds. Large-scale comparisons have compared the behavior of these measures, albeit often only at the sign level, e.g. does changing a setting produce the same a change in the underlying DNN and a generalization measure with the same sign (Jiang et al., 2019a; Dziugaite et al., 2020a).
@@
A notable property of DNNs is that to first order, their generalization performance is robust to modest changes in network size, hyperparameters, optimisers, or stopping criteria. This raises a natural question: are generalization measures themselves equally robust to such changes? If a generalization measure captures the basic underlying factors that drive DNN performance, one would expect it to exhibit a similar degree of robustness.
@@
The main finding of this paper is that many of the post-mortem bounds studied in the literature are fragile: small, ostensibly irrelevant training tweaks that barely affect the underlying DNN, can flip a measure’s shape, trend, or scale. To test these fragilities systematically, we evaluate measures under controlled perturbations that hold data and architecture, fixed while nudging a single training knob. Three stressors structure the audit in the main text: learning‑curve behavior as sample size increases (§3); temporal behavior, including when the model has interpolated (§4); and responses to data complexity via label noise and dataset swaps (§5). In the supplementary material, we show more tests of fragility for a much wider set of measures.
@@
In contrast to the post-mortem bounds, we show that a function-based marginal-likelihood PAC-Bayes bound successfully tracks dataset complexity and learning-curve scaling (Valle-Pérez & Louis, 2020). This bound is based on a simple measure of inductive bias, the probability that a random set of weights produces a function that fits the training data set. Because it operates in function space and does not depend explicitly on training dynamics, this bound is inherently insensitive to the choice of optimiser or most training hyperparameters. This insensitivity is both a strength, making the bound robust rather than fragile, and a limitation, as it prevents the bound from reflecting how training procedures influence generalization, including effects such as feature learning. Nonetheless, the strong performance of this relatively simple approach suggests that developing more sophisticated function-space generalization measures could be a fruitful direction for future research.
@@
Finally, we sketch a few potential causes of fragility, in the hope they can lead to new insights. We show that some of the qualitative failures of norm-based bounds with sample size appear to track a recent prediction of qualitative changes in the scaling of norms in the simpler setting of linear regression (Zhang & Louis, 2025). We also exploit a symmetry in scale‑invariant networks and prove a non‑asymptotic equivalence between fixed learning‑rate with fixed weight decay and a matched run with exponentially increasing learning rate and time‑varying weight decay; the two training procedures compute the same predictor at every iterate (§6; cf. Li & Arora, 2019). Under this invariance lever, magnitude‑sensitive post‑mortem measures can inflate by orders of magnitude while test error remains flat. This suggests that generalization measures should carefully track key invariances present in the underlying DNNs.
The widespread fragility we observe suggests that some generalization measures may be showing correlation in some limited settings that may not be related to causation. Our position is therefore not only that many measures are fragile, but also that investigators should carefully audit their measures for fragility, using the kinds of tests we propose in this paper.
Our contribution:
  • We formulate a compact fragility audit for generalization measures that targets training-hyperparameter stability, post‑interpolation temporal behavior, and data‑complexity response.
  • We provide systematic evidence that popular post‑mortem measures—path, spectral and Frobenius norms, flatness proxies, and deterministic PAC–Bayes surrogates—change their qualitative predictions under mild optimiser or hyperparameter tweaks even when the underlying DNN accuracy is stable.
  • We prove an equivalence for scale‑invariant networks that matches fixed schedules to exponentially increasing learning‑rate schedules with time‑varying weight decay, yielding a controlled invariance lever that isolates function from parameter scale and exposes magnitude sensitivity.
@@ 2 Related work
Early theory framed capacity via VC and Rademacher analyses (Mendelson, 2002; Anthony & Bartlett, 2002; Bartlett & Mendelson, 2001; Koltchinskii, 2001). As overparameterization became the norm, evidence accumulated that uniform-convergence explanations often miss the mark: networks can fit random labels yet generalize on natural data (Zhang et al., 2016), some bounds fail to track risk (Nagarajan & Kolter, 2019), and benign overfitting can arise even at interpolation (Bartlett et al., 2020). The field responded with diagnostics that foreground invariances, algorithmic dependence, and data interactions.
2.1 Capacity-oriented diagnostics: norms, margins, distance from initialization
@@
The first works in this subfield translated classical notions into deep settings. Spectrally-normalized margin bounds tied test error to Lipschitz-like control and margins, making scale explicit (Bartlett et al., 2017). Path- and norm-based views clarified how depth and weight scales shape effective complexity (Neyshabur et al., 2015, 2019). Empirically, modeling the distribution of margins—not just the minimum—improves predictiveness across trained families (Jiang et al., 2018; Novak et al., 2018), and sample complexity can depend on norms rather than width (Golowich et al., 2018). In practice, enforcing Lipschitz continuity often improves out-of-sample performance (Gouk et al., 2020; Yoshida & Miyato, 2017). Distance-from-initialization and movement-from-pretraining offer reference-aware, width-robust complexity surrogates with both empirical support and bounds (Li et al., 2018; Zhou et al., 2021; Neyshabur et al., 2017a), while optimization dynamics link large margins to implicit bias in separable regimes (Soudry et al., 2018).
2.2 Geometry-oriented diagnostics: flatness and sharpness
@@
A parallel line approached generalization through loss-landscape geometry. The “flat minima generalize better” intuition predates deep learning (Hochreiter & Schmidhuber, 1997) and resurfaced when large-batch training was observed to converge to sharper minima with worse test error (Keskar et al., 2017). Because raw sharpness is parameterization-dependent, normalized and magnitude-aware definitions were proposed and shown to correlate more robustly with test error (Dinh et al., 2017; Tsuzuku et al., 2020; Kim et al., 2022; Jang et al., 2022). These ideas also shaped algorithms: SAM explicitly optimizes a local worst-case neighborhood and typically improves accuracy; adaptive variants and ablations probe when and why it helps (Foret et al., 2020; Kwon et al., 2021; Andriushchenko & Flammarion, 2022). Independent probes—weight averaging, landscape visualizations, and training-dynamics analyses—add evidence that broader valleys often accompany better generalization (Izmailov et al., 2018; Li et al., 2017; Cohen et al., 2021). Stochastic optimization theory offers a mechanism: noise scale and heavy-tailed gradient noise can bias learning toward flatter regions (Smith et al., 2017; Jastrzebski et al., 2017; Simsekli et al., 2019; McCandlish et al., 2018).
2.3 Algorithm-aware certificates: PAC-Bayes as operational measures
@@
Partly in response to limits of uniform convergence, PAC-Bayesian analysis made algorithm dependence explicit by trading empirical risk against a posterior–prior divergence (McAllester, 1999; Seeger, 2002; Catoni, 2007; Langford & Shawe-Taylor, 2002). The framework became operational when non-vacuous deep-network certificates were obtained by optimizing stochastic posteriors (Dziugaite & Roy, 2017), or simply working with the GP prior (Valle-Pérez et al., 2018; Zhang et al., 2021). Subsequent work tightened certificates with data-aware priors (Dziugaite, 2018; Pérez‐Ortiz et al., 2020), and with compression-flavored posteriors that reduce effective description length (Arora et al., 2018a; Lotfi et al., 2022). Treating PAC-Bayes itself as a training objective yields models that are both accurate and tightly certified, while extensions broaden the scope to adversarial risk and fast/mixed-rate regimes (Rodríguez-Gálvez et al., 2024). However, one worry is that while these optimised measures may provide quantitatively tighter bounds, they do so at the expense of qualitative insight (Arora et al., 2018a).
Other perspectives.
@@
Complementary lenses provide boundary conditions any credible measure should respect. Algorithmic stability ties generalization to insensitivity under data perturbations (Bousquet & Elisseeff, 2002; Hardt et al., 2015); information-theoretic analyses bound excess risk by the information a learning rule extracts from the sample (Russo & Zou, 2016; Xu & Raginsky, 2017). Description-length and compression viewpoints link compressibility to generalization (Blier & Ollivier, 2018; Taiji Suzuki, 2020). Linearized regimes (NTK/GP) delineate when very wide nets behave like their kernel limits (Jacot et al., 2018; Lee et al., 2020b). Finally, double-descent phenomena and scaling laws offer external checks on diagnostics (Belkin et al., 2019; Kaplan et al., 2020).
Meta-evaluations and synthesis.
@@
Large-scale comparisons underscore that no single candidate explains generalization under all interventions (Jiang et al., 2019b; Dziugaite et al., 2020b), motivating a measures-as-diagnostics mindset. A recent critique shifts the lens from predictiveness to tightness, showing that uniformly tight bounds are out of reach in overparameterized settings and clarifying what such tightness could mean (Gastpar et al., 2023). Our focus is complementary: rather than tightness per se, we study fragility—how otherwise informative measures can fail under innocuous hyperparameter changes that leave underlying DNN performance essentially unchanged.
@@ 3 Training-hyperparameter fragility
The path norm (definition in App. A) is often regarded as the strongest norm-based proxy for modern ReLU networks: it respects layerwise rescalings and has supported some of the sharpest norm-style capacity statements and practical diagnostics (Neyshabur et al., 2015, 2017b; Gonon et al., 2023). Precisely because it is a strong candidate, it is a natural place to look for fragility.
In Fig. 1 we show a triptych for ResNet‑50 on FashionMNIST, nudging exactly one hyperparameter at a time, either the optimizer or the learning rate (no early stopping), while keeping the data, model, and pipeline fixed. Each curve shows test error (black), the raw path norm (blue), and the path norm divided by the empirical margin (orange); axes are logarithmic and error bars reflect seed variability.
@@
In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower. Both drop faster than the underlying ResNet‑50 test error. Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior. A minor adjjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows. These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
The contrasting behavior of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression. In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer. More precisely, with $p\in[1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if $r\leq 2(p-1)$  (Zhang & Louis, 2025). This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm. This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (Zhang & Louis, 2025). For the more complex setup of path-norm for ResNet‑50 on FashinMNIST in Fig. 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
See App. A for precise definitions for the measures we consider as well as further experiments, including the same stress test across Frobenius, margin, spectral, PAC‑Bayes, and VC‑style proxies.

### v3

@@ Abstract
In this position paper, we argue that many post-mortem generalization measures—those computed on trained networks—are fragile: small training modifications that barely affect the performance of the underlying deep neural network can substantially change a measure’s value, trend, or scaling behavior. For example, minor hyperparameter changes, such as learning rate adjustments or switching between SGD variants, can reverse the slope of a learning curve in widely used generalization measures such as the path norm. We also identify subtler forms of fragility. For instance, the PAC-Bayes origin measure is regarded as one of the most reliable, and is indeed less sensitive to hyperparameter tweaks than many other measures. However, it completely fails to capture differences in data complexity across learning curves. This data fragility contrasts with the function-based marginal-likelihood PAC-Bayes bound, which does capture differences in data-complexity, including scaling behavior, in learning curves, but which is not a post-mortem measure. Beyond demonstrating that many post-mortem bounds are fragile, this position paper also argues that developers of new measures should explicitly audit them for fragility.
@@ 1 Introduction to the position statement
Classical generalization theories—based on VC dimension, Rademacher complexity and related tools—have a long history in machine learning. They were developed with two intertwined aims: to furnish tight mathematical guarantees and to explain generalization. For deep networks the same aspiration has driven a profusion of bounds. While tight bounds have proven elusive, with some exceptions (dziugaite2017computing; prezortiz2020tighter; valleprez2020generalization; lotfi2022pacb), the expectation has been that these measures can generate qualitative insight into the generalization behavior of deep neural networks (DNNs). These bounds have typically focused on weights after training, which is why they are sometimes called post-mortem bounds. Large-scale comparisons have compared the behavior of these measures, albeit often only at the sign level, e.g. does changing a setting produce the same change in the underlying DNN and a generalization measure with the same sign (jiang2019fantastic; dziugaite2020in).
@@
A notable property of DNNs is that to first order, generalization performance is robust to modest changes in network size, hyperparameters, optimizers, or stopping criteria. This raises a natural question: do generalization measures share this robustness? If a generalization measure captures the basic underlying factors that drive DNN performance, one would expect it to exhibit a similar degree of robustness. The main position this paper argues for is that: Many of the post-mortem bounds studied in the literature are fragile: small training tweaks that barely affect the underlying DNN performance can drastically change a measure’s behaviour. Investigators should carefully audit their generalization measures for fragility, using the kinds of tests we propose in this paper.
@@
To test these fragilities systematically, we develop a fragility audit that evaluates measures under controlled perturbations that hold data and architecture fixed while nudging a single training knob. Three stressors structure the audit in the main text: learning‑curve behavior as sample size increases (§3); temporal behavior, including when the model has interpolated (§4); and responses to data complexity via label noise and dataset swaps (§5). §6 reveals that the tightest existing bound fails across both its bias and complexity components. In the supplementary material, we test a much wider set of measures and develop an expanded fragility audit.
@@
In contrast to the post-mortem bounds, we show that a function-based marginal-likelihood PAC-Bayes bound successfully tracks dataset complexity and learning-curve scaling (valleprez2020generalization). This bound is based on a simple measure of inductive bias, the probability that a random set of weights produces a function that fits the training data set. Because it operates in function space and does not depend explicitly on training dynamics, this bound is inherently insensitive to the choice of optimizer or most training hyperparameters. This insensitivity is both a strength, making the bound robust rather than fragile, and a limitation, as it prevents the bound from reflecting how training procedures influence generalization, including mechanisms such as feature learning. Nonetheless, the strong performance of this relatively simple approach suggests that developing more sophisticated function-space generalization measures could be a fruitful direction for future research.
@@
Finally, we sketch a few potential causes of fragility, in the hope they can lead to new insights. We show that some of the qualitative failures of norm-based bounds with sample size appear to track a recent prediction of qualitative changes in how norms scale with data set size in the simpler setting of linear regression (zhang2025closed). We also exploit a symmetry in scale‑invariant networks and prove a non‑asymptotic equivalence between fixed learning‑rate with fixed weight decay and a matched run with exponentially increasing learning rate and time‑varying weight decay; the two training procedures compute the same predictor at every iterate (§7; cf. li2019an). Under this invariance lever, magnitude‑sensitive post‑mortem measures can inflate by orders of magnitude while test error remains flat. This suggests that generalization measures should carefully track key invariances present in the underlying DNNs.
We issue a call to action: Further work to understand the causes of fragility is essential. Otherwise it is hard to assess whether a generalization measures reflects a true causal relationship or not.
Our contribution: We (1) formulate a stress-test protocol for hyperparameter stability, temporal drift, and data complexity; (2) demonstrate that popular post-mortem measures are fragile, often reversing trends or failing to track data difficulty under benign tweaks that leave the underlying predictor unchanged.; (3) introduce a quantitative score for measure instability; and (4) prove a non-asymptotic equivalence between fixed and exponential LR schedules for scale-invariant nets, exposing flaws in magnitude-sensitive measures.
@@ 2 Related work
From classical capacity to modern practice. Early theory framed capacity via VC and Rademacher analyses (vapnik1998statistical; anthony1999neural; bartlett2001rademacher; koltchinskii2001rademacher). As overparameterization became the norm, evidence accumulated that uniform-convergence explanations often miss the mark: networks can fit random labels yet generalize on natural data (zhang2016understandingc), some bounds fail to track risk (nagarajan2019uniform), and benign overfitting can arise even at interpolation (bartlett2020benign). The field responded with diagnostics that foreground invariances, algorithmic dependence, and data interactions.
@@
Capacity-oriented diagnostics: norms, margins, distance from initialization. Early work translated classical notions into deep settings. Spectrally-normalized margin bounds tied test error to Lipschitz-like control and margins, making scale explicit (bartlett2017spectrallyb). Path- and norm-based views clarified how depth and weight scales shape effective complexity (neyshabur2015norm; neyshabur2019theb). Empirically, modeling the distribution of margins, not just the minimum, improves predictiveness across trained families (jiang2018predicting; novak2018sensitivity), and sample complexity can depend on norms rather than width (golowich2018size). In practice, enforcing Lipschitz continuity often improves out-of-sample performance (gouk2020regularisation; yoshida2017spectral). Distance-from-initialization and movement-from-pretraining offer reference-aware, width-robust complexity surrogates with both empirical support and bounds (li2018on; zhou2021understanding; neyshabur2017exploring), while optimization dynamics link large margins to implicit bias in separable regimes (soudry2018the).
@@
Geometry-oriented diagnostics: flatness and sharpness. A parallel line approached generalization through loss-landscape geometry. The “flat minima generalize better” intuition predates deep learning (hochreiter1997flatb) and resurfaced when large-batch training was observed to converge to sharper minima with worse test error (keskar2016on). Because raw sharpness is parameterization-dependent, normalized and magnitude-aware definitions were proposed and shown to correlate more robustly with test error (dinh2017sharpb; tsuzuku2020normalized; kim2022scale; jang2022reparametrization). These ideas also shaped algorithms: SAM explicitly optimizes a local worst-case neighborhood and typically improves accuracy; adaptive variants probe when and why it helps (foret2020sharpnessb; kwon2021asam; zela2022towards). Independent probes—weight averaging, landscape visualizations, and training-dynamics analyses—add evidence that broader valleys often accompany better generalization (izmailov2018averagingb; li2017visualizing; cohen2021gradient). Stochastic optimization theory offers a mechanism: noise scale and heavy-tailed gradient noise can bias learning toward flatter regions (smith2017donb; jastrzbski2017three; simsekli2019tail; mccandlish2018an).
@@
Algorithm-aware certificates: PAC-Bayes as operational measures. Partly in response to limits of uniform convergence, PAC-Bayesian analysis made algorithm dependence explicit by trading empirical risk against a posterior–prior divergence (mcallester1999someb; seeger2002pac; catoni2007pac; langford2002pac). The framework became operational when non-vacuous deep-network certificates were obtained by optimizing stochastic posteriors (dziugaite2017computing), or simply working with the GP prior (valleprez2018deepb; zhang2021why). Subsequent work tightened certificates with data-aware priors (dziugaite2018data; prezortiz2020tighter), and with compression-flavored posteriors that reduce effective description length (arora2018strongerc; lotfi2022pacb). Treating PAC-Bayes itself as a training objective yields models that are both accurate and tightly certified, while extensions broaden the scope to adversarial risk and fast/mixed-rate regimes (rodrguezglvez2024mixed). However, one worry is that while these optimised measures may provide quantitatively tighter bounds, they do so at the expense of qualitative insight (arora2018strongerc).
@@
Other perspectives. Complementary lenses provide boundary conditions any credible measure should respect. Algorithmic stability ties generalization to insensitivity under data perturbations (bousquet2002stability; hardt2016train); information-theoretic analyses bound excess risk by the information a learning rule extracts from the sample (russo2016controlling; xu2017information). Description-length and compression viewpoints link compressibility to generalization (blier2018the; suzuki2020compression). Linearized regimes (NTK/GP) delineate when very wide nets behave like their kernel limits (jacot2018neuralb; lee2019wide). Finally, double-descent phenomena and scaling laws offer external checks on diagnostics (belkin2019reconciling; kaplan2020scaling).
@@
Meta-evaluations and synthesis. Large-scale comparisons underscore that no single candidate explains generalization under all interventions (jiang2019fantastic; dziugaite2020in), motivating a measures-as-diagnostics mindset. A recent critique shifts the lens from predictiveness to tightness, showing that uniformly tight bounds are out of reach in overparameterized settings and clarifying what such tightness could mean (gastpar2023fantastic). Our focus is complementary: rather than tightness per se, we study fragility—how otherwise informative measures can fail under innocuous hyperparameter changes that leave underlying DNN performance essentially unchanged.
@@ 3 Training-hyperparameter fragility
The path norm (definition in App. A) is often regarded as the strongest norm-based proxy for modern ReLU networks: it respects layerwise rescalings and has supported some of the sharpest norm-style capacity statements and practical diagnostics (neyshabur2015norm; neyshabur2017a; gonon2023a). Precisely because it is a strong candidate, it is a natural place to look for fragility.
In Fig. 1 we show a triptych for ResNet‑50 on FashionMNIST, nudging exactly one hyperparameter at a time, either the optimizer or the learning rate (no early stopping), while keeping the data, model, and pipeline fixed. Each curve shows test error (black), the raw path norm (blue), and the path norm divided by the empirical margin (orange). In the left plot, we run SGD with momentum at $\mathrm{LR}=0.01$: both path‑norm curves start at a value of roughly ($\sim 10^{5}$) and decrease more or less log‑linearly with increasing sample size, with the margin‑normalized curve tracking about a decade lower. Both drop faster than the underlying ResNet‑50 test error. Changing the optimizer to Adam but keeping $\mathrm{LR}=0.01$ (middle panel) results in much smaller path norms (now near $10^{-1}$) and striking non-monotonic behavior. A minor adjustment of Adam’s learning rate to $0.001$ (right panel)—and the metrics jump back up to $\sim 10^{5}\text{--}10^{6}$ but now collapse by four orders of magnitude as data size grows. These widely different behaviors arise purely from modest hyperparameter tweaks that barely affect the underlying ResNet‑50 test error.
@@
The contrasting behavior of the path-norm measure appears to mirror the simpler, analytically tractable case for linear regression. In that setting, for a fixed $\ell_{r}$ weight norm, qualitatively different scaling of the norm with sample size obtains depending on the $\ell_{p}$‑minimizer. More precisely, with $p\in(1,2]$, the $\ell_{r}$ weight norm is monotonic in $n$ if $r\leq 2(p-1)$ and plateaus if r>2(p1)r>2(p-1) (zhang2025closed). This implies that fixing which norm you measure does not fix its learning‑curve scaling, because the solution selected by optimization—through implicit (or explicit) bias, e.g., favoring different $\ell_{p}$‑minimizers—can induce different sample‑size scalings of that same norm. This change in qualitative behavior was also illustrated in diagonal linear networks where the implicit bias can be tuned by varying the learning dynamics (zhang2025closed). For the more complex setup of path-norm for ResNet‑50 on FashionMNIST in Fig. 1, the optimizer and learning‑rate choices may analogously shift the deep model’s implicit bias, toggling the path‑norm scaling between monotone and U‑shaped, just as was seen for linear regression and deep linear networks.
See App. A for precise definitions for the measures we consider as well as further experiments, including the same stress test across Frobenius, margin, spectral, PAC‑Bayes, and VC‑style proxies [^1] All experiments in this work are fully reproducible with minimum hardware requirements. The anonymized code can be found at https://anonymous.4open.science/r/Prior_bound-8A6C/..